# SMURF: Scalable method for unsupervised reconstruction of flow in 4D flow MRI

### v1

@@ Abstract
Index Terms:
4D flow MRI, Hemodynamic reconstruction, Unsupervised segmentation, Cardiovascular diagnostics.
@@ I Introduction
By capturing time-resolved data, 4D flow MRI facilitates the dynamic volumetric segmentation of vascular structures, enhancing sensitivity to anatomical abnormalities. This segmentation process establishes vascular boundaries essential for computing wall shear stress (WSS), a biomarker that relates to endothelial behavior, vascular remodeling, and aneurysm progression [7, 8, 9]. Additionally, once segmented, these structures support the analysis of intravascular pressure distributions, which estimate pressure gradients and vascular resistance [7]. Kinetic energy losses and turbulence metrics help quantify flow disturbances and assess hemodynamic efficiency. By integrating these metrics, 4D flow MRI provides clinicians insights into vascular function, aiding pathology identification and intervention planning [10, 11].
[…]
@@ II-F Optimizer and parameter initialization
  • Initialize the magnitude mean parameters $\{\mu_{m}\}_{m=1}^{M}$ to be evenly spaced between $-1.0$ and $1.0$. Set the variance parameters $\{\sigma^{2}_{m}\}_{m=1}^{M}$ and the diagonal elements of the velocity covariance matrices $\{\mathbf{\Sigma}_{m}\}_{m=1}^{M}$ to 1.0. The method remains robust to these initial values as long as the data is standardized.
@@ III-A2 In vitro Poiseuille flow
[…]
@@ III-A3 In vivo internal carotid artery aneurysm
[…]
@@ III-B Verification on the synthetic 4D flow
[…]
@@
$T$: Voxel set classified as flow in ground truth. $P$: Voxel set classified as flow in predicted segmentation. $\neg T$: Voxel set classified as background in ground truth. $\neg P$: Voxel set classified as background in predicted segmentation. $\cap$: Intersection of two sets. $\cup$: Union of two sets. $\setminus$: Set difference. ||$\cdot$|| ⋅ |: Cardinality (number of voxels in a set).
@@ IV Conclusion
We test SMURF on synthetic, in vitro, and in vivo data. In synthetic data, SMURF recovers velocities in high-noise conditions and maintains sub-voxel segmentation accuracy, outperforming or matching existing methods. In the in vitro experiment, SMURF aligns reconstructed velocities with theoretical profiles, reduces velocity errors compared to raw measurements, and recovers underlying flow structures. In in vivo data, SMURF isolates anatomical anomalies and decreases velocity divergence residuals, supporting the assessment of patient-specific 4D flow measurement.
@@ Appendix A Neural network architectures for geometry and velocity fields
The four-dimensional input vector, which encapsulates both spatial and temporal dimensions, passes through a Fourier feature embedding before entering the modified MLP. This technique transforms the input vectors into high-frequency signals before processing them through the modified MLP. The Fourier feature mapping function $\gamma:\mathbb{R}^{d}\rightarrow\mathbb{R}^{2m}$ is formulated as $\gamma(\mathbf{x})=[\cos(\mathbf{B}\mathbf{x}),\sin(\mathbf{B}\mathbf{x})]$, where $\mathbf{B}\in\mathbb{R}^{m\times d}$ is composed of entries drawn from a Gaussian distribution $\mathcal{N}(0,\sigma^{2})$ and $\sigma>0$ is a user-specified hyperparameter. In our work, $d=4$ and we set the dimensionality of the high-frequency signal space to $2562m=$ 256, comprising 128 sine and 128 cosine components. The standard deviation parameter is set to $\sigma=5.0$.

### v3

@@ I Introduction
By capturing time-resolved data, 4D flow MRI facilitates the dynamic volumetric segmentation of vascular structures, enhancing sensitivity to anatomical abnormalities. This segmentation process establishes vascular boundaries essential for computing wall shear stress (WSS), a biomarker that relates to endothelial behavior, vascular remodeling, and aneurysm progression [42, 10, 22]. Additionally, once segmented, these structures support the analysis of intravascular pressure distributions, which estimate pressure gradients and vascular resistance [42]. Kinetic energy losses and turbulence metrics help quantify flow disturbances and assess hemodynamic efficiency. By integrating these metrics, 4D flow MRI provides clinicians insights into vascular function, aiding pathology identification and intervention planning [33, 8, 32, 31].
Despite clinical benefits, several challenges limit the broader adoption of 4D flow MRI. In the absence of contrast enhancement, poor and inhomogeneous signals require long acquisition times, causing patient discomfort and reducing clinical throughput. Imaging artifacts degrade data quality and complicate flow analysis. Limited spatial resolution, often around $1\,\text{mm}$ in voxel size, creates uncertainties when defining vascular boundaries and resolving velocities in complex flow regions. Current segmentation and velocity reconstruction methods rely heavily on manual expert annotations and isolated processes, introducing workflow inefficiencies, variability, and errors. These factors directly affect the accuracy of derived hemodynamic parameters. A scalable, unsupervised method that automatically segments vascular geometries and reconstructs velocity fields at higher resolutions can address these problems.
@@ II-F Optimizer and parameter initialization
  • Initialize the magnitude mean parameters $\{\mu_{m}\}_{m=1}^{M}$ to be evenly spaced between 1.0-1.0 and $1.0$. Set the variance parameters $\{\sigma^{2}_{m}\}_{m=1}^{M}$ and the diagonal elements of the velocity covariance matrices $\{\mathbf{\Sigma}_{m}\}_{m=1}^{M}$ to 1.0. The method remains robust to these initial values as long as the data is standardized.
@@ III-A2 In vitro Poiseuille flow
We acquired 4D flow MRI data at Purdue MRI Facility on a GE Discovery MR750 3T scanner. The setup included steady Poiseuille flow at about $30\,\text{cm/s}$ mean velocity through a Polydimethylsiloxane (PDMS) channel of a quarter-inch diameter. A Compuflow 1000 MR piston pump drove the flow. We used a 60/40 water-glycerol mixture with a density of $1110\,\text{kg/m}^{3}$ and a viscosity of $0.0035\,\text{Pa}\cdot\text{s}$.
We acquired data at $1.0\times 1.0\times 1.0\,\text{mm}^{3}$ and apply ZIP $\times$ 2 interpolation to $0.5\times 0.5\times 0.5\,\text{mm}^{3}$. We captured nine timepoints over the cardiac cycle, with a repetition time TR = $9.15\,\text{ms}$, an echo time TE = $3.50\,\text{ms}$, flip angle $8^{\circ}$, venc = $70\,\text{cm/s}$, and total scan time of about 4 minutes.
We also performed a TOF scan with $0.5078\times 0.5078\times 1.0\,\text{mm}^{3}$ resolution, flip angle $8^{\circ}$, TR = $3.552\,\text{ms}$, and TE = $1.252\,\text{ms}$. We segmented the images semi-automatically using ITK-Snap [41]. For comparisons with the 4D flow-segmented geometry, we automatically registered two surfaces using point-cloud registration from Python’s Open3D library [43].
@@ III-A3 In vivo internal carotid artery aneurysm
An ICA aneurysm was imaged at Northwestern Memorial Hospital using a 3 T MRI scanner (Skyra, Siemens Healthcare, Erlangen, Germany). The protocol included TOF angiography at $0.4\times 0.4\times 0.6\,\text{mm}^{3}$ and 4D flow MRI with TE/TR = $2.997/6.4\,\text{ms}$, flip angle = $15^{\circ}$, venc = $80\,\text{cm/s}$, temporal resolution = $44.8\,\text{ms}$, and spatial resolution = $1.09\times 1.09\times 1.30\,\text{mm}^{3}$. For further details on the processing, please refer to [5].
@@ III-B Verification on the synthetic 4D flow
We further assess segmentation quality using Accuracy, F1 Score, and Jaccard Index defined in Table I. Fig. 3(b–d) shows that all methods classify background better than flow, as indicated by $\text{Accuracy}>\text{F1 Score}>\text{Jaccard Index}$. Interestingly, all methods achieve their best performance at 20% noise, likely due to an interplay between the spatial blurring inherent to 4D flow (which tends to overpredict flow voxels) and the erosive effects of noise that offset this overprediction. SMURF maintains stable performance up to 50% noise, consistently achieving scores above 90% across all metrics, with the ranking trend remaining $\text{SMURF}>\text{PCD}>\text{SDM}$. At 60% noise, SMURF underpredicts flow, resulting in a decline in performance.
@@
$T$: Voxel set classified as flow in ground truth. $P$: Voxel set classified as flow in predicted segmentation. $\neg T$: Voxel set classified as background in ground truth. $\neg P$: Voxel set classified as background in predicted segmentation. $\cap$: Intersection of two sets. $\cup$: Union of two sets. $\setminus$: Set difference. $|\cdot|$: Cardinality (number of voxels in a set).
@@ IV Conclusion
We test SMURF on synthetic, in vitro, and in vivo data. In synthetic data, SMURF recovers velocities in high-noise conditions and maintains sub-voxel segmentation accuracy, outperforming or matching existing methods. In the in vitro experiment, SMURF aligns reconstructed velocities with theoretical profiles, reduces velocity errors compared to raw measurements, and recovers underlying flow structures. In in vivo data, SMURF isolates anatomical anomalies and decreases velocity divergence residuals, supporting the assessment of patient-specific 4D flow measurement. Future work will extend flow validation by benchmarking against independent in vitro flow references, e.g., volumetric particle tracking velocimetry with stochastic tracking [15, 16, 30].
@@ Appendix A Neural network architectures for geometry and velocity fields
The four-dimensional input vector, which encapsulates both spatial and temporal dimensions, passes through a Fourier feature embedding before entering the modified MLP. This technique transforms the input vectors into high-frequency signals before processing them through the modified MLP. The Fourier feature mapping function $\gamma:\mathbb{R}^{d}\rightarrow\mathbb{R}^{2m}$ is formulated as $\gamma(\mathbf{x})=[\cos(\mathbf{B}\mathbf{x}),\sin(\mathbf{B}\mathbf{x})]$, where $\mathbf{B}\in\mathbb{R}^{m\times d}$ is composed of entries drawn from a Gaussian distribution $\mathcal{N}(0,\sigma^{2})$ and $\sigma>0$ is a user-specified hyperparameter. In our work, $d=4$ and we set the dimensionality of the high-frequency signal space to $2m=256$, comprising 128 sine and 128 cosine components. The standard deviation parameter is set to $\sigma=5.0$.